OpenStudy (anonymous):

Math each geometric sequence with the correct sum. a1 = 343, an = -1, r = -1/7

OpenStudy (texaschic101):

sn = (a1 - r^n) / 1 - r a1 = 343 n = -1 r = -1/7 now lets sub s-1 = (343 - (-1/7^-1)/1 - (-1/7) s-1 = (343 + 1/7^-1)/ 1 + 1/7 s-1 = (343 + .1429^-1) / 7/7 + 1/7 s-1 = (343 + 7) / (8/7) s-1 = 340/(8/7) s-1 = 340 * 7/8 s-1 = 2380/8 s-1 = 297.50 you should probably get a second opinion on this